OpenStudy (anonymous):

Solve the equation: 11q^2-44=0

OpenStudy (phi):

what do you get if you divide both sides (all terms) by 11 ?

OpenStudy (anonymous):

q^2-4 my teacher taught me how to factor stuff like 16s^2+8s+1 (which is (8s+2)(8s-1)) but I don't know what the =0 part has to do with anything.

OpenStudy (phi):

you have an equation. Often, you use factoring to solve equations to solve this equation, I would (as a first step) divide both sides by 11 or if you like, factor an 11 out of each term on the left side

OpenStudy (anonymous):

would I have \[11(q^2-4) = 0\] ?

OpenStudy (phi):

yes. notice you have a difference of squares, so you can factor q^2 - 4 use the rule \[ (a^2 - b^2) =(a+b)(a-b) \]

OpenStudy (anonymous):

quadratic?

OpenStudy (anonymous):

I've never seen that formula. And yes kind of

OpenStudy (anonymous):

Because wouldn't factoring 11 out of both sides mean I would have = 0/11

OpenStudy (phi):

difference of squares arises from (a+b)(a-b) (if you multiply it out, you get a^2 - b^2 )

OpenStudy (anonymous):

Oh so you mean how there's no b because it ended up being 0?

OpenStudy (phi):

** Because wouldn't factoring 11 out of both sides mean I would have = 0/11 *** we "do the same thing" to both sides of an equation, but that does NOT include factoring. You can think of factoring as re-writing ... it does not change the value ... just how it looks. In other words if you have 2*1+2*2 = 6 you would re-write the left side by factoring out a 2: 2*(1+2) = 6 notice it is still equal. You do not have to also factor the 6

OpenStudy (anonymous):

Okay so its basically the same as factoring things without the =0

OpenStudy (phi):

yes, but in this case, we will actually "solve" the equation for values of q that make the equation true. so far you have \[ 11(q^2-4)= 0 \] now factor q^2 - 4

OpenStudy (anonymous):

\[11(q-2)(q+2) = 0\]

OpenStudy (anonymous):

do u understand?

OpenStudy (phi):

now the last part. It is an interesting (but maybe not obvious) idea that if you have A*B=0 A times B is 0 then either A is zero or B is zero (or maybe both A and B are zero) we use that idea to solve \[ 11 (q-2) (q+2) = 0 \]

OpenStudy (anonymous):

I've never seen that and I'll have to go upstairs to watch that video give me a minute

OpenStudy (anonymous):

@ilikequadraticsokay

OpenStudy (phi):

you can watch the video when you have time, it's for future reference.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

@phi @ilikequadraticsokay

OpenStudy (anonymous):

here is the working out and answer

OpenStudy (anonymous):

I appreciate that picture but it's not the quadratic formula

OpenStudy (anonymous):

Can you just show me how to do it? SOrry if I sound pushy, it's just i'm running out of time:/

OpenStudy (anonymous):

I understand that difference of squares, just not how to solve for q

OpenStudy (solomonzelman):

11q^2-44=0 factor out of 11 11(q^2-4)=0 divide both sides by 11. q^2-4=0 add four to both sides, and you can do it from there.

OpenStudy (solomonzelman):

@ilikequadraticsokay, good?

OpenStudy (anonymous):

q^2=4 q=2 ??????

OpenStudy (solomonzelman):

yes, but it is ±2.

OpenStudy (anonymous):

What is that symbol? I haven't seen it yet. Do you think y teacher would be okay with me putting =2? She hasn't taught us this material yet, just how to factor....

OpenStudy (solomonzelman):

it means (plus minus) that basically says that number negative or positive. In this case it means that the answer is both, negative 2, and positive 2.

OpenStudy (anonymous):

Ah, that's in reference to parabolas right?

OpenStudy (solomonzelman):

If your teacher taught you how to factor, I think you should know this. You can plug in both, it doesn't matter b/c\[(-a)^2=a^2\] and so it is with 2\[(-2)^2=2^2\]

OpenStudy (solomonzelman):

you will get the exact same result if you plug in '2' or '-2' if the teacher asks you to check, then plug in both and state that each works.

OpenStudy (anonymous):

How do i solve for x in (4x + 1)(4 - 1)=0?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yes....

OpenStudy (anonymous):

How do i solve for x in (4x + 1)(4 - 1)=0?

OpenStudy (solomonzelman):

if you see that\[a \times b=0\] it would be true that either a=0 or b=0 @ilikequadraticsokay, right?

OpenStudy (anonymous):

I guess?

OpenStudy (solomonzelman):

here you have a similar thing\[(4x+1)\times (4-1)=0\] you know that 4-1 is 3 (and of course not zero.) \[(4x+1)3=0\] divide both sides by 3

OpenStudy (anonymous):

i meant to type 4x-1 not 4-1

OpenStudy (solomonzelman):

so \[ (4x+1)(4x-1)=0\] so either 4x+1=0 or 4x-1=0 WHY? remember what I told you about a times b = zero I am sure you can now do it!

OpenStudy (anonymous):

\[x \pm 1/4\] ?

OpenStudy (solomonzelman):

when you say x±1/4 that would mean x+1/4 or x-1/4 You menat x= ±1/4 Either way this is incorrect Solve each one 4x+1=0 4x-1=0

OpenStudy (anonymous):

No? 4x+1=0 and 4x-1=0, 4x=-1 and 4x=1, both sides divided by 4 are x==1/4 and x=1/4

OpenStudy (solomonzelman):

I was trying to confuse you, sorry.... I do that sometimes to make sure the helper understood. I see you did! You were correct x= ±1/4

OpenStudy (anonymous):

Hahaha okay... thanks!

OpenStudy (anonymous):

so if i have the variable inside like 11(q^2-4)=0 how do i go about that?

OpenStudy (solomonzelman):

you can divide both sides by 11.

OpenStudy (anonymous):

????? can you do that out? would it just be q^2-4)=0? sorry i keep asking for someone to do these out, i have like twenty other problems like these to do....

OpenStudy (solomonzelman):

yes, it would be just q^2-4)=0 b/c  see?